can you please help me with this problem
the expression 8m+5 represents the total cost in dollars, including shipping, for a certain number of music CD's, m. based on the expression .what is the total coast for 4 CDs

Answer by stanbon(75887)

the expression 8m+5 represents the total cost in dollars, including shipping, for a certain number of music CD's, m. based on the expression .what is the total coast for 4 CDs
-------
Since 8m+5 is the cost of "m" CD,
the cost of 4 CDs is 8*4+5 = 32+5 = $37.00
====================
Cheers,
Stan H.
====================